OpenStudy (kingj):

What is the sum of the arithmetic sequence 8, 14, 20 …, if there are 24 terms?

OpenStudy (anonymous):

S = 2a + (n-1)d Where a = first number in sequence, n is nth term you are looking for and d is the common difference.

OpenStudy (anonymous):

sorry, it should be: S = n/2( 2a + (n-1)d)

OpenStudy (mertsj):

\[s _{n}=\frac{n}{2}(a _{1}+a _{n})\] To find the 24th term use: \[a _{n}=a _{1}+(n-1)d=8+(24-1)6=146\] \[s _{n}=\frac{24}{2}(6+146)\] \[s _{n}=1824\]

OpenStudy (mertsj):

There. I think I finally got it right.

OpenStudy (kingj):

THANK YOU

OpenStudy (mertsj):

yw

OpenStudy (anonymous):